# Network-Optimised Spiking Neural Network (NOS) Scheduling for 6G O-RAN: Spectral Margin and Delay-Tail Control

### v1

@@ V Experimental Setup and Results
A single coupling gain $g$ is fixed once at the worst case (largest $\rho(W)$, $\Delta=20$ ms) by tail-parity calibration: NOS $p_{99.9}$ is matched to PF $p_{99.9}$ at that design point (the match_p999_floor rule in code). The same $g$ is then used for all topologies and delays. In numerics the spike slope is absorbed, so $H=1$. For scheduler $s\in\{\mathrm{NOS},\mathrm{PF},\mathrm{BP}\}$ the working margin follows Definition 1 with $\delta^{(s)}=k_{\star}(\Delta_{\mathrm{eff}}^{(s)})-\mathbf{1}_{\{s=\mathrm{NOS}\}}\,g\,\rho(W)$, and we take $\eta_{\mathrm{PF}}=\eta_{\mathrm{BP}}=1$.
Utilisation is mapped from $\delta$ by the saturating law implemented in the code, $\mathrm{AUC}=\gamma_{\mathrm{eff}}\;\frac{\kappa_{u}\,\delta}{1+\kappa_{u}\,\delta},$ with $\kappa_{u}=25$ and a clamp at $0.05\,\gamma_{\mathrm{eff}}$ when $\delta\leq 0$. Tail quantities use Corollary 2 with $\mu_{\min}=0.12$ pkts/slot and $\bar{x}=\min\{1,\mathrm{AUC}/\gamma\}$, yielding $p_{99.9}=\frac{1}{\mu_{\min}\bar{x}}\sqrt{\frac{\log 1000}{\kappa_{\theta}\,\delta}}\quad\text{and}\quad\mathrm{MaxQ}\approx\sqrt{\frac{\log 1000}{\kappa_{\theta}\,\delta}},$ with $\kappa_{\theta}=4.0$. Two normalisations are reported: the nominal-cap mode $\bar{x}=\mathrm{AUC}/\gamma_{c}^{\mathrm{nom}}$ used for default tables and averaged curves, and the effective-cap mode $\bar{x}=\mathrm{AUC}/\gamma_{\mathrm{eff}}$ shown in figures as a sensitivity check.
Figure 2: AUC across pair2, line4, and ring8 at $\Delta\in\{5,12,20\}$ ms. Nominal-cap normalisation; setup as in Sec. V.

### v2

@@ V Experimental Setup and Results
A single coupling gain $g$ is fixed once at the worst case (largest $\rho(W)$, $\Delta=20$ ms) by tail-parity calibration: NOS $p_{99.9}$ is matched to PF $p_{99.9}$ at that design point (the match_p999_floor rule in code). The same $g$ is then used for all topologies and delays. In numerics the spike slope is absorbed, so $H=1$. For scheduler $s\in\{\mathrm{NOS},\mathrm{PF},\mathrm{BP}\}$ the working margin follows Definition 1 with $\delta^{(s)}=k_{\star}(\Delta_{\mathrm{eff}}^{(s)})-\mathbf{1}_{\{s=\mathrm{NOS}\}}\,g\,\rho(W)$, and we take $\eta_{\mathrm{PF}}=\eta_{\mathrm{BP}}=1$. To avoid confusion with classifier ROC-AUC, we report a normalised utilisation headroom $\mathrm{AUC}_{\mathrm{util}}$ over the load range, defined as the area under the utilisation–load curve divided by the effective cap. In code this is implemented by the saturating mapping below, and the plotted quantity is the dimensionless ratio $\mathrm{AUC}_{\mathrm{util}}/\gamma_{\mathrm{eff}}\in[0,1]$. Values in the 0.40.40.70.7 range therefore indicate moderate headroom under the chosen margins and delays, not poor forecasting accuracy. Utilisation is mapped from $\delta$ by the saturating law implemented in the code, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{AUC}_{\mathrm{util}}=\gamma_{\mathrm{eff}}\;\frac{\kappa_{u}\,\delta}{1+\kappa_{u}\,\delta},$} with $\kappa_{u}=25$ and a clamp at $0.05\,\gamma_{\mathrm{eff}}$ when $\delta\leq 0$.
Tail quantities use Corollary 2 with $\mu_{\min}=0.12$ pkts/slot and $\bar{x}=\min\{1,\mathrm{AUC}_{\mathrm{util}}/\gamma\}$, yielding $p_{99.9}=\frac{1}{\mu_{\min}\bar{x}}\sqrt{\frac{\log 1000}{\kappa_{\theta}\,\delta}}\quad\text{and}\quad\mathrm{MaxQ}\approx\sqrt{\frac{\log 1000}{\kappa_{\theta}\,\delta}},$ with $\kappa_{\theta}=4.0$. Two normalisations are reported: the nominal-cap mode $\bar{x}=\mathrm{AUC}_{\mathrm{util}}/\gamma_{c}^{\mathrm{nom}}$ used for default tables and averaged curves, and the effective-cap mode $\bar{x}=\mathrm{AUC}_{\mathrm{util}}/\gamma_{\mathrm{eff}}$ shown in figures as a sensitivity check.
Figure 2: Utilisation headroom AUC across pair2, line4, and ring8 at $\Delta\in\{5,12,20\}$ ms. Nominal-cap normalisation; setup as in Sec. V.